A fence of length 100 feet is used to enclose a rectangular dog pen. The pen is 31 feet longer than it is wide. What are the dimensions of the pen?

Let x-width
31x+31x+x+x

Answer by TimothyLamb(4379)

x = length
y = width
---
2L + 2w = 100
L = w + 31
---
put the system of linear equations into standard form
---
2x + 2y = 100
x - y = 31
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = length = 40.5 feet
y = width = 9.5 feet
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php